# NeurIPS 2025 Workshop on Time Series for Health

### v1

@@ Abstract
Reinforcement learning (RL) applications in Clinical Decision Support Systems (CDSS) frequently encounter skepticism from practitioners regarding inoperable dosing decisions. We address this challenge with an end-to-end approach for learning optimal drug dosing and control policies for dual vasopressor administration in intensive care unit (ICU) patients with septic shock. For realistic drug dosing, we apply action space design that accommodates discrete, continuous, and directional dosing strategies in a system that combines offline conservative Q-learning with a novel recurrent modeling in a replay buffer to capture temporal dependencies in ICU time-series data. Our comparative analysis of norepinephrine dosing strategies across different action space formulations reveals that the designed action spaces improve interpretability and facilitate clinical adoption while preserving efficacy. Empirical results [^1] Code for this paper is publicly available at https://github.com/wzoustanford/ucsf_rl/ on eICU and MIMIC demonstrate that action space design profoundly influences learned behavioral policies. The proposed methods achieve improved patient outcomes of over 15% in survival improvement probability, while aligning with established clinical protocols.
@@ 1 Introduction
Reinforcement learning recommendations are promising to serve as clinician support, but when a seasoned clinician encounters a proposed dosing adjustment as small as $0.01$ mg, the model’s credibility erodes. This disconnect illuminates a need to bridge the gap between theoretically optimal RL policies and recommendations that are clinically actionable, interpretable, and trusted.
Clinical interventions in Intensive Care Units (ICU’s) are inherently heterogeneous, spanning discrete choices and continuous adjustments jayaraman2024primer ; eghbali2021patient ; qiu2022latent ; tosca2024model . This diversity necessitates action space formulations that can flexibly capture the full spectrum of clinical practice while remaining tractable for RL algorithms. Without careful action space design, RL policies risk generating unrealistic or impractical recommendations that clinicians are unlikely to adopt. More specifically, in a Clinical Decision Support System (CDSS) hunt1998effects ; garg2005effects for septic shock, vasopressor dosing requires dynamic control across multiple phases—initiation, titration, and weaning. Current clinical practices adhere to international standards of applying norepinephrine as the first line vasopressor and vasopressin as the second-line vasopressor. While reinforcement learning offers a natural framework for this sequential decision-making problem, existing work kalimouttou2025optimal ; komorowski2018artificial has predominantly focused on single vasopressor control and early treatment phases, overlooking the reality of titration during the full length of therapy.
To address these challenges, we investigate the patterns of both vasopressors, especially the impact of norepinephrine dosing patterns with effective action space design for clinical drug dosing towards a realistic CDSS system. Concretely, we adopt the offline settings of the Deep Q-learning (DQN) mnih2015human with recurrent replay kapturowski2018recurrent , and conservative Q-learning (CQL) kumar2020conservative . Integrating them together as a potential CDSS system, we validate the action space design for dosing in treatment trajectories.
We find that discrete and directional-discrete action spaces are not only more interpretable and aligned with clinician expectations, but also improve survival probability by over 15% compared to continuous dosing. Their sparsity and higher dimensionality, combined with recurrent and replay frameworks on short sequences, enable more generalizable models. These results show the potential of deep reinforcement learning and action-space design for building effective CDSS systems.
@@ 2 Prior Work
For action space design in RL, earlier work addressed factored tavakoli2018action , hierarchical kulkarni2016hierarchical , continuous action spaces tessler2019action ; delalleau2019discrete kapturowski2018recurrent addressed the importance of recurrent sequence learning in replays for application in Atari games, and  schrittwieser2020mastering progressed recurrent learning into model-based RL applications. Offline RL methods such as conservative Q-learning kumar2020conservative , implicit Q-learning kostrikov2021offline and advantage weighted Actor-Critic nair2020awac applied different regularization and policy constraints. Model-based offline RL approaches kidambi2020morel focused on modeling transition dynamics. In the clinical literature, eghbali2021patient applied RL algorithm to sedation dosing in ICUs. A collection of dosing techniques was given in  tosca2024model , focusing on referencing cancer and chemotherapy treatments.  qiu2022latent applied a latent constrained batch algorithm to address the risks of incorrect dosing heparin in ICU. However, the use of RL to systematically design and evaluate vasopressor dosing strategies for sepsis shock remains under-explored.
@@ 3 Algorithms
We propose an offline Q-learning framework with recurrent experience replay, focusing on effective dosing action-spaces with conservatism objective from conservative Q-learning.
3.1 Vasopressor Control Problem Formulation
We formulate the vasopressor control problem with the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space of vasopressor decisions, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward, and $\gamma\in[0,1]$ is the discount factor. The state $s_{t}\in\mathcal{S}\subseteq\mathbb{R}^{d}$ comprises patient information, physiological measurements [^2] The measurements include mean blood pressure (MBP), lactate levels, blood urea nitrogen, serum creatinine, (total) fluid, urine of the hour, corticosteroid and Sequential Organ Failure Assessment (SOFA) score. and organ support [^3] The support includes mechanical ventilation, renal replacement therapy.. Action space $\mathcal{A}$ contains a binary decision for vasopressin ($vp_{1}$) and dosing decisions for norepinephrine ($vp_{2}$). The dosing decisions are formulated in Section 3.3. The reward function contains two objectives: maximizing the chance of survival and keeping vasopressors in dosage limit. Given a dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\}$ of historical ICU trajectories collected under behavioral policy $\pi_{b}$ (clinical practice), our goal is to learn an optimal policy $\pi^{*}$ that maximizes expected cumulative reward.
3.2 Q-Learning and conservative Q-Learning (CQL)
Standard Q-learning seeks to learn the action-value function $Q(s,a)$ representing the expected cumulative reward from taking action $a$ in state $s$ and following the optimal policy thereafter. The Q-function satisfies the Bellman optimality: $Q^{*}(s,a)=\mathcal{R}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})]$. In the offline setting, Q-learning minimizes the temporal difference (TD) mean-squared error:
$\mathcal{L}_{\text{TD}}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{\theta}(s,a)-(r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(s^{\prime},a^{\prime}))\right)^{2}\right],$
where $Q_{\theta}$ is the learned Q-network with parameters $\theta$, and $Q_{\theta^{-}}$ is a target network updated periodically for stability. We apply a version of double Q-learning [^4] We use two Q nets, taking Q-value as $q=\text{min}(q_{1},q_{2})$; two target nets, taking target Q as $q^{\prime}=\text{min}(q_{1}^{\prime},q_{2}^{\prime})$. Standard Q-learning in offline settings may have over-estimation bias due to distributional shift, causing over-optimistic Q-value estimates.
As an improvement to Q-learning, CQL kumar2020conservative makes a correction by adding a conservative regularization term that penalizes Q-values for out-of-distribution (OOD) actions:
$\mathcal{L}_{\text{CQL}}(\theta)=\alpha\cdot\mathbb{E}_{s\sim\mathcal{D}}\left[\log\sum_{a}\exp(Q_{\theta}(s,a))-\mathbb{E}_{a\sim\mathcal{D}(a|s)}[Q_{\theta}(s,a)]\right]+\mathcal{L}_{\text{TD}}(\theta),$
where $\alpha$ controls the strength of the conservative penalty. When $\alpha=0$, it reduces to Q-learning.
3.3 Effective Dosing Action Space Design
We propose an action-space design to integrate clinical implementable practice with reinforcement learning. While continuous dosing variables are elegant, they may not agree with clinical realities of established protocols with discrete thresholds. Additionally, continuous action spaces compound RL research challenges due to exploration complexity and sparse rewards. Our interdisciplinary approach addresses this gap through three complementary action-space formulations: continuous dosage (for comparison), clinically meaningful discrete blocks, and directional stepwise adjustments. The action space $\mathcal{A}$={($vp_{1}$$vp_{2}$)} encompasses dual vasopressor dosing, enabling both computational tractability and clinical feasibility.
@@ 3.3 Effective Dosing Action Space Design
Continuous dosing (baseline). We use a continuous dosing variable for vasopressor norepinephrine allowing it to freely move move in a range. The action space is $\mathcal{A}=\{(vp_{1b},vp_{2c})\}$, where $vp_{1b}\in\{0,1\}$ is a binary variable, and $vp_{2c}\in[0,0.5]$ (mcg/kg/min).
@@
Block discrete dosing. We define the dosing action to be the discrete values in the same range as the continuous dosing. The discrete values only vary across meaningful discrete dosing blocks. We explore different numbers of blocks to validate the impact of effective dosing. The action space is $\mathcal{A}=\{(vp_{1b},vp_{2d})\}$, and $vp_{2c}\in\text{range}(\delta/2,0.5,\delta)$ (mcg/kg/min) where $\delta$ is the minimum step [^5] Due to international clinical practice, we ensure the dosing is above zero for the first-line vasopressor..
@@
Stepwise directional dosing. We define the the action space at each time step to be a directional and stepwise change. $\mathcal{A}=\{(vp_{1b},vp_{2s})\}$, and $vp_{2s}$ is a stepwise change from the current dose, for example, $vp_{2s}\in[{-0.1,-0.05,0,+0.05,+0.1}]$ (mcg/kg/min). The state space is extended to include a one-hot vector that stores and keeps track of the current block-discrete dosage of norepinephrine.
@@ 3.4 Recurrent Experience Replay with Effective Dosing
Our approach adapts the recurrent experience replay algorithm kapturowski2018recurrent to address the temporal dependencies in ICU vasopressor control. The algorithm employs a prioritized sequential buffer $\mathcal{RB}=\{\mathcal{S}_{e}\}$, storing treatment episodes $\mathcal{S}_{e}=\{(s_{t},a_{t},r_{t},s_{t+1})_{t\in e}\}$, to capture time-sensitive patterns of patient responses to vasopressor adjustments. In contrast to the sample-wise Q-function application in traditional Q-learning, we leverage an LSTM network to model the sequential nature of ICU states and interventions, recognizing that vasopressor effects unfold over time. The sum-tree data structure prioritizes episodes with larger temporal difference (TD) errors schaul2015prioritized , enabling learning on challenging clinical scenarios. This combination of recurrent processing and specialized action-space design addresses the ICU’s demands for both immediate responsiveness and patient stability.
@@ 4 Experiments
We apply the algorithms on the eICU pollard2018eicu , and MIMIC-IV johnson2023mimic datasets. The data is preprocessed with $16$ features and is split with ratios $70/15/15$ for train/val/test. In our RL formulation, the $vp_{1}$ (vasopressin) dosing is set as a binary variable because the clinical practice uses a single max dosage. We primarily experiment with the titration and weaning of $vp_{2}$ (norepinephrine)  [^6] See Appendix A and for detailed experiments. B. See Appendix C and D for model implementation and architecture. See Appendix E for hyperparameters and Appendix F for FQE off-policy evaluation.. The reward definition of the reinforcement learning model is summarized in Table 1.
Table 1: Definition of reward in the RL model.
Rule
Mortality +10 if the patient survived, and -10 otherwise
Vaso. penalty Each time step, aggregate to reward: $0.1\cdot(vp_{1}+vp_{2}\cdot 2-1.0))$
Our work compares several action-space formulations for vasopressor control in Table 2. We evaluate Q improvement ($\Delta Q$)—the gain in model-recommended actions over clinician actions—averaged across time steps. Each Q-function maps state–action pairs to a scalar value of expected future rewards, making $\Delta Q$ directly comparable across models. The Binary $vp_{1}$ model treats vasopressin as a binary variable. he Dual Mixed model represents $vp_{2}$ as a continuous variable in the range $[0,0.5]$ (mcg/kg/min). The Dual Block Discrete (BD) model discretizes $vp_{2}$ into $3$, 55, or $10$ bins. The Dual Stepwise model defines $vp_{2s}\in[{-\text{max\_step},...,-0.05,0,+0.05,...,+\text{max\_step}}]$ (mcg/kg/min). We report concordance (C.) between model recommendations and clinician actions. For $vp_{1}$, concordance is binary agreement; for $vp_{2}$, concordance holds when the recommended dose falls within the defined bin edges. Finally, the LSTM recurrent experience replay uses $vp_{2}$ discretized into 10 bins, trained with sequence length 5 and a 2-step burn-in. See model implemetation details in Appendix C.
Table 2: Comparison of RL models including Binary $vp_{1}$, Dual Mixed, Dual Block Discrete, Dual Stepwise models with the best conservatism levels ($\alpha$) selected with the validation set.

### v2

@@ Abstract
Reinforcement learning (RL) applications in Clinical Decision Support Systems (CDSS) frequently encounter skepticism because models may recommend inoperable dosing decisions. We propose an end-to-end offline RL framework for dual vasopressor administration in Intensive Care Units (ICUs) that directly addresses this challenge through principled action space design. Our method integrates discrete, continuous, and directional dosing strategies with conservative Q-learning and incorporates a novel recurrent modeling using a replay buffer to capture temporal dependencies in ICU time-series data. Our comparative analysis of norepinephrine dosing strategies across different action space formulations reveals that the designed action spaces improve interpretability and facilitate clinical adoption while preserving efficacy. Empirical results on eICU and MIMIC demonstrate that action space design profoundly influences learned behavioral policies. Compared with baselines, the proposed methods achieve more than $3\times$ expected reward improvements, while aligning with established clinical protocols  [^1] Code is publicly available at https://github.com/wzoustanford/vaso_rl/.
@@ 1 Introduction
Clinical interventions in ICUs are inherently heterogeneous, spanning discrete choices and continuous adjustments jayaraman2024primer ; eghbali2021patient ; qiu2022latent ; tosca2024model . This diversity places strong demands on action space in Reinforcement Learning (RL): they should capture the full spectrum of clinical practice while remaining tractable for learning. Without careful action space design, RL policies risk generating unrealistic or impractical recommendations that clinicians are unlikely to adopt. Specifically, in Clinical Decision Support System (CDSS) hunt1998effects ; garg2005effects for septic shock, clinicians apply norepinephrine (first-line vasopressor) and vasopressin (second-line vasopressor) over multiple phases—initiation, titration, and weaning. While reinforcement learning offers a natural framework for this sequential decision-making problem, prior work kalimouttou2025optimal ; komorowski2018artificial often focuses on single vasopressor control and early treatment phases, with fewer examining how titration evolves throughout the entire treatment period.
In this work, we study the effectiveness of action space design in influencing the learned dosing policies. Concretely, we adopt the offline settings of the Deep Q-learning (DQN) mnih2015human with recurrent replay kapturowski2018recurrent to capture temporal dependencies, and incorporate conservative Q-learning (CQL) kumar2020conservative to ensure policy safety. Integrating them together as a potential CDSS system, we validate the action space design for dosing in treatment trajectories. Using this integrated offline RL framework as a CDSS, we evaluate discrete, continuous, and directional dosing action spaces and analyze their impact on both norepinephrine and vasopressin control over the full treatment horizon.
Our results show that discrete and directional-discrete action spaces are not only more interpretable, aligned with clinician expectations, but they also significantly improve model performance compared to continuous dosing. Their sparsity and higher dimensionality, combined with the recurrent replay framework enable more generalizable models that achieve more than $3\times$ expected reward improvements with weighted importance sampling offline-policy evaluation. These results show the potential of deep reinforcement learning and action space design for building effective CDSS systems.
@@ 2 Prior Work
For action space design in RL, earlier work addressed factored tavakoli2018action , hierarchical kulkarni2016hierarchical , continuous action spaces tessler2019action ; delalleau2019discrete kapturowski2018recurrent addressed the importance of recurrent sequence learning in replays for application in Atari games, and  schrittwieser2020mastering progressed recurrent learning into model-based RL applications. Offline RL methods such as conservative Q-learning kumar2020conservative , implicit Q-learning kostrikov2021offline and advantage weighted Actor-Critic nair2020awac applied different regularization and policy constraints. Model-based offline RL approaches kidambi2020morel focused on modeling transition dynamics. In the clinical literature, eghbali2021patient applied RL algorithm to sedation dosing in ICUs. tosca2024model provided acollection of dosing techniques focusing on referencing cancer and chemotherapy treatments.  qiu2022latent applied a latent constrained batch algorithm to address the risks of incorrect dosing heparin in ICU. However, the use of RL to systematically design and evaluate vasopressor dosing strategies for sepsis shock remains under-explored.
@@ 3 Algorithms
We propose an offline Q-learning framework that employs dosing action space design with a conservative Q-learning objective, and incorporates a recurrent replay buffer to model ICU time-series.
3.1 Dosing Action Space Design
We introduce an action space design that bridges the gap between reinforcement learning optimization and clinical feasibility. The action space contains a binary decision for vasopressin ($vp_{1}$) and dosing decisions for norepinephrine ($vp_{2}$) which takes values in the range $(0,0.5]$ (mcg/kg/min) [^2] Due to international clinical practice, we ensure the dosing is above zero for the first-line vasopressor.. Specifically, three action space designs for $vp_{2}$ are proposed: continuous dosage, clinically-aligned discrete intervals based on standard protocols, and directional dose adjustments that mirror clinical decision-making. Our action space $\mathcal{A}$={($vp_{1}$$vp_{2}$)} models dual vasopressor administration, maintaining model performance while preserving the tractability of real-world intensive care interventions. This design enables deployment within existing clinical workflows while maintaining the expressiveness required for effective policy learning.
@@ 3.1 Dosing Action Space Design
Continuous dosing. A continuous dosing variable is applied for vasopressor norepinephrine allowing it to freely move in a range. The action space is $\mathcal{A}=\{(vp_{1},vp_{2})\}$, where $vp_{1}\in\{0,1\}$ is a binary variable, and $vp_{2}\in(0,0.5]$ (mcg/kg/min).
@@
Block discrete dosing. We define the dosing action to be the discrete values in the same range as in the continuous dosing. The discrete values only vary across meaningful discrete dosing blocks. We explore different numbers of blocks to validate the impact of effective dosing. The action space is $\mathcal{A}=\{(vp_{1},vp_{2})\}$, where $vp_{2}\in\text{range}(\delta/2,0.5,\delta)$ (mcg/kg/min) and $\delta$ is bin size computed from the number of bins.
@@
Stepwise directional dosing. We define the the action space at each time step to be a directional and stepwise change. The action space is $\mathcal{A}=\{(vp_{1},vp_{2})\}$, where $vp_{2}$ is a stepwise change from the current dose. For example, $vp_{2}\in\{-0.1,-0.05,0,+0.05,+0.1\}$ (mcg/kg/min). The state space is extended to include a one-hot vector that stores and keeps track of the current discrete dosage of norepinephrine.
3.2 Q-Learning for Vasopressor Control
We formulate the vasopressor control problem with $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space of vasopressor decisions, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward, and $\gamma\in[0,1]$ is the discount factor. The state $s_{t}\in\mathcal{S}\subseteq\mathbb{R}^{d}$ comprises patient information, physiological measurements  [^3] The measurements include mean blood pressure (MBP), lactate levels, blood urea nitrogen, serum creatinine, (total) fluid, urine of the hour, corticosteroid, and Sequential Organ Failure Assessment (SOFA) score. , and organ support  [^4] The support includes mechanical ventilation and renal replacement therapy.. The action space $\mathcal{A}$ contains a binary decision for vasopressin ($vp_{1}$) and dosing decisions for norepinephrine ($vp_{2}$). The dosing decisions are formulated in Section 3.1. The reward function contains two objectives: maximizing the chance of survival and keeping vasopressors in dosage limit. Given a dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\}$ of historical ICU trajectories collected under behavioral policy $\pi_{b}$ (clinical practice), our goal is to learn an optimal policy $\pi^{*}$ that maximizes expected cumulative reward.
Q-Learning. Q-learning seeks to learn the action-value function $Q(s,a)$ representing the expected cumulative reward by taking action $a$ in state $s$ and following the optimal policy thereafter. The Q-function satisfies the Bellman optimality: $Q^{*}(s,a)=\mathcal{R}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})]$. In the offline setting, Q-learning minimizes the temporal difference (TD) mean-squared error:
$\mathcal{L}_{\text{TD}}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{\theta}(s,a)-(r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(s^{\prime},a^{\prime}))\right)^{2}\right],$ (1)
where $Q_{\theta}$ is the learned Q-network with parameters $\theta$, and $Q_{\theta^{-}}$ is a target network updated periodically for stability. We apply a version of double Q-learning van2016deep with two Q networks and two corresponding target networks.
As an improvement to Q-learning, we optionally add a conservative term kumar2020conservative (conservative Q-Learning) to penalize Q-values for out-of-distribution actions, with $\alpha$ as strength of the penalty:
$\mathcal{L}_{\text{CQL}}(\theta)=\alpha\cdot\mathbb{E}_{s\sim\mathcal{D}}\left[\log\sum_{a}\exp(Q_{\theta}(s,a))-\mathbb{E}_{a\sim\mathcal{D}(a|s)}[Q_{\theta}(s,a)]\right]+\mathcal{L}_{\text{TD}}(\theta).$
@@ 3.3 Recurrent Experience Replay with Effective Dosing
Our approach applies the recurrent experience replay algorithm kapturowski2018recurrent to address the temporal dependencies in ICU vasopressor control. The algorithm employs a prioritized sequential buffer $\mathcal{RB}=\{\mathcal{S}_{e}\}$, storing treatment episodes $\mathcal{S}_{e}=\{(s_{t},a_{t},r_{t},s_{t+1})_{t\in e}\}$, to capture time-sensitive patterns of patient responses to vasopressor adjustments. In contrast to the sample-wise Q-function application in traditional Q-learning, we leverage an LSTM network to model the sequential nature of ICU states and interventions, recognizing that vasopressor effects unfold over time. In addition, we incorporate the sum-tree data structure to prioritize episodes with larger temporal difference (TD) errors schaul2015prioritized , improving learning in challenging clinical scenarios. The recurrent model combines with specialized action space design to improve reinforcement learning for patient stability.
@@ 4 Experiments
We conduct multiple experiments demonstrating that our action space designs are both learnable and effective. We first validate learnability and model behavior through pilot evaluations with $\Delta$ Q-value and Fitted Q-Evaluation, using a simplified reward (Section 4.1). We then show concrete gains through full off-policy evaluation under a comprehensive clinical reward (Section 4.2).
4.1 Learnability of Clinically Aligned Action Spaces
We evaluate our dosing action space design with the reward function defined in Table 1. The objective is to evaluate whether each action space formulation supports meaningful value learning by tracking Q-value improvements and estimating policy quality using Fitted Q-Evaluation (FQE) on sepsis treatment data from eICU pollard2018eicu and MIMIC-IV johnson2023mimic datasets  [^5] See Appendix A for more data and processing details..
Type Rule of the reward function
Mortality: +10+10 if the patient survived, and 10-10 otherwise
Vaso. penalty: At each time step, aggregate to reward: $0.1\cdot(vp_{1}+vp_{2}\cdot 2-1.0))$
Table 1: Definition of the simple reward function.
Action Spaces. Following Section 3.1, we compare five formulation of action space: (i) the Binary $vp_{1}$ model, which treats vasopressin as a binary variable; (ii) the Dual Mixed model, which presents $vp_{2}$ as a continuous variable in the range $(0,0.5]$ (mcg/kg/min); (iii) the Dual Block Discrete (Dual BD) model, which discretizes $vp_{2}$ into $3$, 55, or $10$ bins; (iv) the Dual Stepwise model, which encodes directional dose adjustments $vp_{2s}\in[{-\text{max\_step},...,-0.05,0,+0.05,...,+\text{max\_step}}]$ (mcg/kg/min); (v) the LSTM Block Discrete (LSTM BD), which discretizes $vp_{2}$ into $10$ bins trained with sequence length $5$ and a $2$-step burn-in.
RL Model Implementation. The RL model leverages a fully-connected network as a Q-function which takes the combined state-action as input. For double-Q learning, during training and inference, Q-value is taken as the minimum of the target outputs, $q=\text{min}(q_{1},q_{2})$. The Q networks are periodically updated by adding weighted parameters of the target networks. The model is trained with the TD error objective (Equation 1)  [^6] See Appendix F and G for model implementation and architecture. See Appendix H for hyper-parameters..
$\Delta$ Q Evaluation. As shown in Table 2, action space designs are evaluated with Q improvement per transition ($\Delta\text{Q /step}$), where $\Delta Q$ is the gain of model-recommended actions over clinician actions. We also measure concordance (C.) between model recommendations and clinician actions. For $vp_{1}$, concordance is binary agreement; for $vp_{2}$, concordance holds when the recommended dosage falls within the defined bin edges.
Table 2: Comparison of RL models. We include Binary $vp_{1}$, Dual Mixed, Dual BD, Dual Stepwise, and LSTM BD models with the best conservatism levels ($\alpha$) selected with the validation set.